Version of May 17, 2007

Tue 22  Wed 23 Thu 24  Fri 25  Sat 26  Sun 27
9.00  Kegel  Stöhr  Hall  Thomas  Passman Meierfrankenfeld
10.00 P. Smith Olshanskii Borovik Grigorchuk  Lashki  Rowley
11.00  coffee
11.30 Leinen Brown Pragacz Stichtenoth Shumyatsky H. Smith
12.30  lunch
14.30 Kuzucuoğlu  Lewis  Shalev  Zelmanov
15.30  Bryant  Zalesski  Roquette  Pride
16.00  Papistas  Makarenko
16.30  coffee  coffee
17.00  Amberg  ∥ talks  ∥ talks  ∥ talks
17.30 Behravesh
18.00 reception
18.30  ∗

Moderators:

9.00 Ali Nesin  P. Smith Bilhan Shumyatsky  Hall  Bryant
14.30  Rowley  Sertöz  Stichtenoth  Zalesski
17.00 A  Asar  Stöhr  H. Smith  Güloğlu
17.00 B  Ünlü  Topuzoğlu  Alizade

∗ Friday, after dinner: memories of Brian Hartley, with his daughter, C. Barber-Brown

Parallel talks

17.00  17.30  18.00  18.30
Wed A  Yalçın  Yalçınkaya  †
B  Alizade  Pusat-Yılmaz Crispin Q.  Kaptanoğlu
Fri A  Asar  Suleymanova Diasamidze
B Çeşmelioğlu  Khan  Yıldız
Sat A  Levchuk  Ercan  ‡
B  Basarab  Güçlükan  Altunbulak Dzhumadil’daev

† Johnson, Arıkan, Torstensson, Russo ‡ Tolstykh, Deloro, Pastori, Sica